# Harnessing quantum back-action for time-series processing

### v2

@@ Abstract
Quantum measurements affect the state of the observed systems via back-action. While projective measurements extract maximal classical information, they drastically alter the system. In contrast, indirect measurements balance information extraction with the degree of disturbance. Considering the prevalent use of projective measurements in quantum computing and communication protocols, the potential benefits of indirect measurements in these fields remain largely unexplored. In this work, we demonstrate that incorporating indirect measurements into a quantum machine-learning protocol known as quantum reservoir computing provides advantages in both execution time scaling and overall performance. We analyze different measurement settings by varying the measurement strength across two benchmarking tasks. Our results reveal that carefully optimizing both the reservoir Hamiltonian parameters and the measurement strength can significantly improve the quantum reservoir computing algorithm performance. Furthermore, our approach demonstrates improved memory performance when compared with state-of-the-art classical feedback protocols. This work provides a comprehensive and practical recipe to promote the implementation of indirect measurement-based protocols in quantum reservoir computing. Moreover, our findings motivate further exploration of experimental protocols that leverage the back-action effects of indirect measurements.
@@ I Introduction
The degree of disturbance induced by a measurement depends on its nature. Conventional projective measurements cause substantial disruption to the quantum system, forcing it into the eigenstate (or a degenerate subspace) associated with the measured outcome. This strong interaction collapses superposition and fully determines the system’s state. In contrast, indirect measurements provide a gentler alternative by using indirect interaction between the system and the measurement apparatus [10.1093/acprof:oso/9780199213900.001.0001]. The strength of the coupling and the duration of this interaction govern both the disturbance introduced by the measurement and the amount of information extracted. Consequently, indirect measurements can provide only partial information about the system, inducing a back-action that depends on the degree of measurement strength. This approach allows the quantum state to retain some coherence, enabling ongoing observation of the system without fully collapsing it. This makes indirect measurements a valuable tool for quantum information processing [PhysRevA.62.012307, SINGH2014141]. Indirect measurements have been both demonstrated and widely used across various quantum platforms [pan2020weak, spagnolo2022experimental, murch2013observing]. Superconducting qubits, for example, are typically measured through indirect measurements using ancillary microwave pulses [naghiloo2019introduction, weber2016quantum]. By adjusting the strength and duration of these pulses, experimenters can tune the measurement strength from weak to strong, with the latter approaching projective measurements in the limit [murch2013observing, Murch2016, doi:10.1126/science.1226897]. In general, the capacity to implement indirect measurements relies on the ability to prepare convenient ancillas [Yasuda2023] or to extract information through projectively measuring only a portion of the system [hu2024overcoming, viqueira2023density].
@@
In this work, we demonstrate that adjusting the strength of quantum measurements enables the optimization of quantum back-action introduced into the system’s evolution, thereby enhancing the performance of machine learning tasks beyond what is achievable without this control. We illustrate our findings within the context of time-series processing using quantum reservoir computing (QRC) [mujal2021opportunities]. QRC belongs to the broader category of quantum machine learning, a field that emerged from the desire to combine machine learning techniques with the unique advantages of quantum systems [dawid2023]. First introduced in [fujii2017harnessing], it is a supervised quantum machine learning technique that leverages the advantages of quantum systems as the exponential growth of their state space with increasing system size. At the same time, it incorporates the simple and efficient training strategy characteristic of classical reservoir computing [jaeger2004harnessing], where the internal dynamics of the reservoir are fixed and only the output layer is trained. Recently, there has been growing interest in QRC and related approaches, with progress both in theoretical advancements [Nakajima2019boosting, Ghosh2019, Ghosh2019prl, Kutvonen2020, martinez2020information, Ghosh2020reconstructing, Tran2020higherorder, Nokkala2021, martinez2021dynamical, Ghosh2021neuromorphic, mujal2021analytical, govia2022nonlinear, Angelatos2021qstatemeas, Nokkala2022, Llodra2022, Domingo2022, Yelin2022, Nokkala2023, GarciaBeni2023, martinez2023finitedim, Krisnanda2023tomographic, mujal2023time, Innocenti2023, Wudarski2023, GarciaBeni2024squeezing, sannia2024dissipation, Nokkala2024, Jaurigue2024memory, Li2024estimating, Kobayashi2024, Domingo2024, Kobayashi2024eco, palacios2024role, Llodra2024] and experimental implementations on diverse platforms [negoro2018machine, Chen2020, Suzuki2022, spagnolo2022experimental, Pfeffer2022, Kubota2023, Yasuda2023, hu2024overcoming, Paternostro2024, Yelin2024].
@@
Interestingly, in various contexts, controlling the degree of decoherence introduced into the reservoir has been shown to improve its memory capacity [martinez2023finitedim, sannia2024dissipation, monzani2024leveraging, palacios2024role]. In this work, we build on this insight by systematically exploring the entire range of indirect measurement regimes and jointly optimizing the measurement strength with the reservoir Hamiltonian parameters, in order to clarify its fundamental role in the reservoir’s performance. By leveraging the ability to tune the measurement strength, we optimize the tradeoff between information extraction and back-action. This allows for an online processing scheme that benefits from linear scaling without suffering drawbacks from decoherence. Instead, by carefully controlling the measurement strength, we can enhance the reservoir’s capacity. To achieve this, we introduce a new figure of merit to compare the performance of the indirect measurement-based protocol, where back-action is controlled, with projective measurement-based protocols that rely on restarting strategies to avoid back-action. By optimizing this metric, we identify the optimal measurement strength that maximizes performance. We illustrate this improvement across different realizations of reservoirs operating in distinct dynamical regimes, validated through two benchmark tasks: the forecasting of the Santa Fe chaotic series [PhysRevA.40.6354] and the short-term memory task . Furthermore, we extend our analysis to a realistic finite resource scenario and show that the enhancement effect is amplified by the favourable scaling properties of the indirect measurement protocol. Finally, we compare our approach with a classical feedback strategy, demonstrating the advantage of exploiting a pure quantum effect such as back-action with respect to incoherent feedback.
@@ II.1 Quantum reservoir computing via indirect measurements
Reservoir computing was proposed as a training-efficient alternative to more complicated recurrent neural network models like Long Short-Term Memory for processing temporal data and time series [jaeger2004harnessing]. Quantum reservoir computing extends the concept by leveraging quantum properties and dynamics, using a quantum system as the reservoir [fujii2017harnessing]. The general pipeline of QRC algorithms is structured in two steps: first the reservoir has to process the time series of interest and then a linear regression model is trained to perform predictions (Fig. 1 (a)). While the latter is a straightforward process, the former includes three stages to be repeated for each element of the dataset: input encoding, time evolution and information extraction (Fig. 1 (b)). At the encoding stage, the information of the $k$-th element of the time series is injected into the amplitudes of the state of a fixed spin of the reservoir [fujii2017harnessing, martinez2021dynamical, mujal2021analytical]. The information is then spread through the system by evolving it for a time $\Delta t$ under a transverse field Ising Hamiltonian $\hat{H}$. Ultimately, we measure a set of observables $\{\langle\hat{O}_{i}\rangle\}$ to extract information from the quantum system. However, when performing projective measurements, it is convenient to reset the system and reprocess the previous $k$ elements without applying measurements before processing the next input, $k+1$. This is because projective measurements erase the stored information in quantum coherences through a destructive process. This procedure, known as the restarting protocol (RSP), has long served as the canonical benchmark in the field, since it introduces no additional disturbance to the system. More recently, several alternative protocols have been proposed that focus on offering more practical implementations while still providing good performances [settino, hu2024overcoming, Kobayashi2024, Chen2020], such as the approach based on indirect measurements introduced in Ref. [mujal2023time]. In this scheme, referred to as the online protocol (OLP), only a small amount of information is lost due to the back-action effect (Fig. 1 (c)), allowing the QRC algorithm to run continuously without restarting at each step in the time series processing. The OLP introduces back-action during the evolution as a tradeoff, allowing for better algorithm scaling by avoiding the need to restart at each input.
@@ III.2 Details of the algorithm pipeline
where $\odot$ indicates the element-wise product. Actually, the CPTP map in this last equation is equivalent to a dephasing channel [martinez2023finitedim]. We can then extract the expectation value of the observable $\hat{\sigma}_{i}^{z}$ on the i-th spin as ${\rm Tr}(\hat{\rho}_{k}\hat{\sigma}_{i}^{z})$. To measure single qubit observables in the $x$ and $y$ direction it is necessary to apply proper rotations to ${\mathcal{L}}_{k}(\hat{\rho}_{k-1})$ before applying $\hat{M}$. For two-spin observables the same description applies. Since the RSP is equivalent to dynamics without back-action, it can be conveniently simulated by setting $g=0$ in this model, for which $M$ reduces to an all-ones matrix and has no effect on the state.
@@ IV.1 Performance analysis in the infinite resources scenario
Previous work [mujal2023time] revealed that the same performance achieved with the RSP can, in some cases, also be obtained with the OLP, despite the introduction of back-action effects. On the basis of this, we analyze the performance of QRC with the OLP at different measurement strengths. Moreover, we consider different external field strengths, $h$, i.e. different dynamical regimes of the reservoir [martinez2021dynamical]. Both quantities are expressed as dimensionless throughout this work.
To quantify the performance of one protocol with respect to the other, we introduce the performance ratio, $P_{R}$, which measures how the QRC algorithm running with the OLP compares with the analogous QRC algorithm with the RSP. Specifically, $P_{R}(g,h)$ is defined as the ratio between the OLP’s sum capacity, given a fixed $g$ and $h$, and the best sum capacity, $h$ wise, obtained with the RSP:
$P_{R}(g,h)=\frac{C_{\Sigma}^{\text{OLP}}(g,h)}{\text{max}_{h^{\prime}}(C_{\Sigma}^{\text{RSP}}(h^{\prime}))}$ (7)
@@
We choose the best-performing RSP rather than the corresponding RSP performance for the same $h$, as we aim for a metric that indicates not only if the OLP performs as well as the RSP but also if the QRC algorithm performs well in general. While this is a strict metric, it avoids focusing on trivial scenarios where both protocols perform poorly. With this definition, a $P_{R}>1$ indicates a practical advantage of the OLP over the RSP. One could also choose to maximize $P_{R}$ separately for each direction, thereby obtaining optimal parameters independently along each axis. In Appendix B, we show that in our case this does not improve the results; however, it could guide an informed choice of the preferred measurement direction when resources are limited or bring advantages if one cares about a specific task with small $\eta$. We average the $P_{R}$ over 50 different reservoir realizations, where the couplings in between the spins are randomly selected each time. In the subsequent plots, the lines indicate the mean values, while the shaded regions denote one standard deviation around the mean.
In Figure 2 we show how $P_{R}$ changes at different $g$ and $h$ values for the forward prediction task on the Santa Fe dataset (top panel), and for the short-term memory task (bottom panel). Each colormap is made up of 33 by 40 points and is computed by assuming that we have access to an infinite amount of resources, i.e. an infinite number of shots for each measured observable. The $P_{R}$ at $g=0$ is set to be 0 since it corresponds to the case in which the measurement apparatus is not interacting with the reservoir and the measurement output is then a random number uncorrelated to the system.
@@
Figure 2: Performance ratio analysis for the forward prediction task (top) and for the short-term memory task (bottom). A $P_{R}$ in the range $[0,1)$ shows that the best performing RSP is performing better than the OLP in that point of the ghg-h map. A $P_{R}$ around 1 shows that the OLP is performing as good as the best performing RSP, and over 1 indicates an over-performance. The optimal parameters and corresponding $P_{R}$ values for each task are: forward prediction — RSP: $h^{\prime*}=0.346$; OLP: $g^{*}=0.355$, $h^{*}=0.273$, $P_{R}=1.006$; short-term memory — RSP: $h^{\prime*}=0.084$; OLP $g^{*}=0.256$ , $h^{*}=0.066$ , $P_{R}=1.029$.
For both tasks, we can highlight the clear emergence of a region for which the OLP achieves comparable performances to the best-performing RSP. In the case of the forward prediction task, this region is spread along the $h$ axis while for the short-term memory task is peaked for $h$ close to $10^{-1}$. This indicates that the back-action effect, which would normally be expected to degrade the performance of QRC, instead provides a positive contribution that compensates for the limited information extracted from the system, enabling the OLP to match the performance of the RSP.
@@
Hence, the emergence of these white and almost-red regions highlights that an indirect measurement-based protocol, if properly tuned, loses nothing in terms of performance compared to the RSP. To stress this last point, in Figure 3, we compare the different sum capacities of the QRC running with the RSP and with the OLP in the case in which $g$ is fixed to a non-optimal value and in the case in which it is fixed to the value that maximized $P_{R}$ in Figure 2. As a first observation, we can refer to [martinez2021dynamical] to justify the dependence of the performance on the $h$ value. In fact, along the horizontal axis, the quantum system undergoes a transition from a localized regime to an ergodic one and it was showcased that being in the ergodic regime is a preferential condition to perform reservoir computing. Moreover, what emerges clearly is that optimizing the measurement strength, rather than simply minimizing back-action, is crucial for improving the protocol.
@@
Figure 3: Sum capacity comparison of the RSP with respect to the OLP with and without optimizing the strength of the measurement $g$. With $g^{*}$ we denote the $g$ value that maximizes the $P_{R}$. The top panel shows the results for the forward prediction task while the bottom panel is for the short-term memory task.
@@
The previous analysis is related to the sum capacity of the QRC across two different tasks. However, for different sub-tasks, defined byăvarying $\eta$, the performance may systematically differ. In this context, Figure 4 compares two points in the ghg-h parameter space explored in Fig. 2, the one that maximized $P_{R}$ and a non-optimal point. We then evaluate the performance for each different sub-task relative to the performance of the RSP at optimal $h$.

### v3

# Present address: ]Qilimanjaro Quantum Tech, Carrer de Veneçuela 74, 08019 Barcelona, Spain

@@ Abstract
Quantum measurements affect the state of the observed systems via back-action. While projective measurements extract maximal classical information, they drastically alter the system. In contrast, indirect measurements balance information extraction with the degree of disturbance. Considering the prevalent use of projective measurements in quantum computing and communication protocols, the potential benefits of indirect measurements in these fields remain largely unexplored. In this work, we show that measurement back-action, a purely quantum effect, can be harnessed within the quantum reservoir computing framework to enhance temporal data processing. Incorporating indirect measurements leads to improved execution-time scaling and overall performance. Our results reveal that carefully optimizing the measurement strength can significantly improve the quantum reservoir computing algorithm performance. Furthermore, our approach demonstrates improved memory performance when compared with state-of-the-art classical feedback protocols. This work provides a comprehensive and practical recipe to promote the implementation of indirect measurement-based protocols in quantum reservoir computing. Moreover, our findings motivate further exploration of experimental protocols that leverage the back-action effects of indirect measurements.
@@ I Introduction
The degree of disturbance induced by a measurement depends on its nature. Conventional projective measurements cause substantial disruption to the quantum system, forcing it into the eigenstate (or a degenerate subspace) associated with the measured outcome. This strong interaction collapses superposition and fully determines the system’s state. In contrast, indirect measurements provide a gentler alternative by using indirect interaction between the system and the measurement apparatus [10.1093/acprof:oso/9780199213900.001.0001]. The strength of the coupling and the duration of this interaction govern both the disturbance introduced by the measurement and the amount of information extracted. Consequently, indirect measurements can provide only partial information about the system, inducing a back-action that depends on the degree of measurement strength. This approach allows the quantum state to retain some coherence, enabling ongoing observation of the system without fully collapsing it. This makes indirect measurements a valuable tool for quantum information processing [PhysRevA.62.012307, SINGH2014141, mujal2023time]. Indirect measurements have been both demonstrated and widely used across various quantum platforms [pan2020weak, spagnolo2022experimental, murch2013observing]. Superconducting qubits, for example, are typically measured through indirect measurements using ancillary microwave pulses [naghiloo2019introduction, weber2016quantum]. By adjusting the strength and duration of these pulses, experimenters can tune the measurement strength from weak to strong, with the latter approaching projective measurements in the limit [murch2013observing, Murch2016, doi:10.1126/science.1226897]. In general, the capacity to implement indirect measurements relies on the ability to prepare convenient ancillas [Yasuda2023] or to extract information through projectively measuring only a portion of the system [hu2024overcoming, viqueira2023density].
@@
We illustrate our findings within the context of time-series processing using quantum reservoir computing (QRC) [mujal2021opportunities]. QRC belongs to the broader category of quantum machine learning, a field that emerged from the desire to combine machine learning techniques with the unique advantages of quantum systems [dawid2023]. First introduced in [fujii2017harnessing], it is a supervised quantum machine learning technique that leverages the advantages of quantum systems as the exponential growth of their state space with increasing system size. At the same time, it incorporates the simple and efficient training strategy characteristic of classical reservoir computing [jaeger2004harnessing], where the internal dynamics of the reservoir are fixed and only the output layer is trained. Recently, there has been growing interest in QRC and related approaches, with progress both in theoretical advancements [Nakajima2019boosting, Ghosh2019, Ghosh2019prl, Kutvonen2020, martinez2020information, Ghosh2020reconstructing, Tran2020higherorder, Nokkala2021, martinez2021dynamical, Ghosh2021neuromorphic, mujal2021analytical, govia2022nonlinear, Angelatos2021qstatemeas, Nokkala2022, Llodra2022, Domingo2022, Yelin2022, Nokkala2023, GarciaBeni2023, martinez2023finitedim, Krisnanda2023tomographic, mujal2023time, Innocenti2023, Wudarski2023, GarciaBeni2024squeezing, sannia2024dissipation, Nokkala2024, Jaurigue2024memory, Li2024estimating, Kobayashi2024, Domingo2024, Kobayashi2024eco, palacios2024role, Llodra2024] and experimental implementations on diverse platforms [negoro2018machine, Chen2020, Suzuki2022, spagnolo2022experimental, Pfeffer2022, Kubota2023, Yasuda2023, hu2024overcoming, Paternostro2024, Yelin2024].
@@
In this work, we reveal that adjusting the strength of quantum measurements enables the optimization of quantum back-action introduced into the system’s evolution, thereby enhancing the performance of machine learning tasks beyond what is achievable without this control. Interestingly, in various contexts, controlling the degree of decoherence introduced into the reservoir has been shown to improve its memory capacity [martinez2023finitedim, sannia2024dissipation, monzani2024leveraging, palacios2024role]. In this work, we build on this insight by systematically exploring the entire range of indirect measurement regimes and jointly optimizing the measurement strength with the reservoir Hamiltonian parameters, in order to clarify its fundamental role in the reservoir’s performance. By leveraging the ability to tune the measurement strength, we optimize the tradeoff between information extraction and back-action. This allows for an online processing scheme that benefits from linear scaling without suffering drawbacks from decoherence. Instead, by carefully controlling the measurement strength, we can enhance the reservoir’s capacity. To achieve this, we introduce a new figure of merit to compare the performance of the indirect measurement-based protocol, where back-action is controlled, with projective measurement-based protocols that rely on restarting strategies to avoid back-action. By optimizing this metric, we identify the optimal measurement strength that maximizes performance. We illustrate this improvement across different realizations of reservoirs operating in distinct dynamical regimes, validated through two benchmark tasks: the forecasting of the Santa Fe chaotic series [PhysRevA.40.6354] and the short-term memory task . Furthermore, we extend our analysis to a realistic finite resource scenario and show that the enhancement effect is amplified by the favourable scaling properties of the indirect measurement protocol. Finally, we compare our approach with a classical feedback strategy, demonstrating the advantage of exploiting a pure quantum effect such as back-action with respect to incoherent feedback.
@@ II.1 Quantum reservoir computing via indirect measurements
Reservoir computing was proposed as a training-efficient alternative to more complicated recurrent neural network models like Long Short-Term Memory for processing temporal data and time series [jaeger2004harnessing]. Quantum reservoir computing extends the concept by leveraging quantum properties and dynamics, using a quantum system as the reservoir [fujii2017harnessing]. The general pipeline of QRC algorithms is structured in two steps: first the reservoir has to process the time series of interest and then a linear regression model is trained to perform predictions (Fig. 1 (a)). While the latter is a straightforward process, the former includes three stages to be repeated for each element of the dataset: input encoding, time evolution and information extraction (Fig. 1 (b)). At the encoding stage, the information of the $k$-th element of the time series is injected into the amplitudes of the state of a fixed spin of the reservoir [fujii2017harnessing, martinez2021dynamical, mujal2021analytical]. The information is then spread through the system by evolving it for a time $\Delta t$ under a transverse field Ising Hamiltonian $\hat{H}$. Ultimately, we measure a set of observables $\{\langle\hat{O}_{i}\rangle\}$ to extract information from the quantum system. However, when performing projective measurements, it is convenient to reset the system and reprocess the previous $k$ elements without applying measurements before processing the next input, $k+1$. This is because projective measurements erase the stored information in quantum coherences through a destructive process. This procedure, known as the restarting protocol (RSP), has long served as the canonical benchmark in the field, since it introduces no additional disturbance to the system.
Beyond full restarting, an alternative strategy is the rewinding protocol (RWP) [Chen2020, mujal2023time]. Instead of repeating the entire past evolution after every projective measurement, the RWP leverages the fading-memory property of the reservoir and rewinds only the last $\tau_{\mathrm{wo}}$ steps. Concretely, at each output extraction the system is re-prepared at time $t_{k}-\tau_{\mathrm{wo}}$ and only the corresponding $N_{\mathrm{wo}}=\tau_{\mathrm{wo}}/\Delta t$ inputs are re-injected before measurement (see Fig. 1(c) in Ref. [mujal2023time]). This reduces the repetition overhead from quadratic to linear in the dataset size while still preventing the accumulation of measurement back-action. Still, the RWP relies on projective measurements and explicit replay of past inputs, making it intrinsically non-online. This has motivated alternative approaches that preserve single forward-pass reservoir evolution while achieving competitive performance [settino, hu2024overcoming, Kobayashi2024], including indirect measurement–based schemes introduced in Ref. [mujal2023time]. In this scheme, referred to as the online protocol (OLP), only a small amount of information is lost due to the back-action effect (Fig. 1 (c)), allowing the QRC algorithm to run continuously without restarting at each step in the time series processing. The OLP introduces back-action during the evolution as a tradeoff, allowing for better algorithm scaling by avoiding the need to restart at each input.
@@ III.2 Details of the algorithm pipeline
where $\odot$ indicates the element-wise product. Actually, the CPTP map in this last equation is equivalent to a dephasing channel [martinez2023finitedim]. We can then extract the expectation value of the observable $\hat{\sigma}_{i}^{z}$ on the i-th spin as ${\rm Tr}(\hat{\rho}_{k}\hat{\sigma}_{i}^{z})$. To measure single qubit observables in the $x$ and $y$ direction it is necessary to apply proper rotations to ${\mathcal{L}}_{k}(\hat{\rho}_{k-1})$ before applying $\hat{M}$. For two-spin observables the same description applies. Since the RSP is equivalent to dynamics without back-action, it can be conveniently simulated by setting $g=0$ in this model, for which $M$ reduces to an all-ones matrix and has no effect on the state. For our purposes, the RWP is simulated identically to the RSP. Because rewinding reproduces the same unperturbed reservoir dynamics after the short washout transient, the ideal RWP coincides with the RSP at the level of state evolution. Their difference manifests only when accounting for finite resources, where the two protocols incur different repetition overheads, but this affects only the practical resource cost and not the simulated dynamics themselves.
@@ IV.1 Performance analysis in the infinite resources scenario
Previous work [mujal2023time] revealed that the same performance achieved by protocols following the unperturbed reservoir dynamics, such as the RSP and the RWP, can, in some cases, also be obtained with the OLP, despite the introduction of back-action effects. On the basis of this, we analyze the performance of QRC with the OLP at different measurement strengths. Moreover, we consider different external field strengths, $h$, i.e. different dynamical regimes of the reservoir [martinez2021dynamical]. Both quantities are expressed as dimensionless throughout this work.
To quantify the performance of the OLP with respect to protocols that preserve the unperturbed reservoir dynamics, we introduce the performance ratio $P_{R}$. Since the RSP and the RWP are equivalent in the ideal limit, as both reproduce the same unperturbed trajectory, we use the RSP as representative of the unperturbed dynamics. In its simplest form, $P_{R}$ compares the OLP at fixed $(g,h)$ with the unperturbed dynamics evaluated at the same field value $h$:
$P_{R}(g,h)=\frac{C_{\Sigma}^{\text{OLP}}(g,h)}{C_{\Sigma}^{\text{RSP}}(h)}.$ (7)
This is the key quantity for our analysis, as it directly reveals whether measurement back-action helps or hinders the performance of the reservoir for that particular configuration.
For completeness, we also consider a stricter performance indicator, $P_{R}^{\min}$, where the denominator is replaced by the best-performing unperturbed reservoir across all scanned values of $h$:
$P_{R}^{\min}(g,h)=\frac{C_{\Sigma}^{\text{OLP}}(g,h)}{\max_{h^{\prime}}\,C_{\Sigma}^{\text{RSP}}(h^{\prime})}.$ (8)
@@
While $P_{R}$ captures the role of back-action for the given reservoir, $P_{R}^{\min}$ answers the stronger question of whether the OLP can outperform even the optimal unperturbed configuration. Because the benchmark in the denominator is the best reservoir available, this metric is intentionally strict: values $P_{R}^{\min}>1$ are attained only when the reservoir is intrinsically very good and, in addition, back-action provides a genuine performance enhancement. One could also choose to maximize $P_{R}$ or $P_{R}^{\min}$ separately for each direction, thereby obtaining optimal parameters independently along each axis. In Appendix B, we show that in our case this does not improve the results; however, it could guide an informed choice of the preferred measurement direction when resources are limited or bring advantages if one cares about a specific task with small $\eta$.
We average the $P_{R}$ over 50 different reservoir realizations, where the couplings in between the spins are randomly selected each time. In the subsequent plots, the lines indicate the mean values, while the shaded regions denote one standard deviation around the mean.
@@
Figure 2: Performance ratio analysis for the forward prediction task (top) and for the short-term memory task (bottom). A $P_{R}$ in the range $[0,1)$ shows that the unperturbed dynamics protocol is performing better than the OLP in that point of the gghh map. A $P_{R}$ around 1 shows that the OLP is performing as good as the RSP, and over 1 indicates an over-performance due to the effect of back-action. The maximum $P_{R}$ and corresponding parameter values for each task are: forward prediction: $g^{*}=0.355$, $h^{*}=0.01$, $P_{R}=2.178$; short-term memory: $g^{*}=0.486$, $h^{*}=0.01$, $P_{R}=27.302$.
In Figure 2 we show how $P_{R}$ varies across different values of $g$ and $h$ for the forward prediction task on the Santa Fe dataset (top panel), and for the short-term memory task (bottom panel). Each colormap is composed of 33 by 40 points and is computed under the assumption of having access to an infinite number of shots for each measured observable. Regions where $P_{R}>1$ indicate that back-action provides an advantage over the corresponding unperturbed reservoir. The $P_{R}$ at $g=0$ is set to be 0 since it corresponds to the case in which the measurement apparatus is not interacting with the reservoir and the measurement output is then a random number uncorrelated to the system.
A consistent and evident performance enhancement is observed across broad regions of the parameter space. In particular, back-action is most beneficial when the unperturbed dynamics struggles the most, such as at very small values of $h$, where the reservoir becomes fully localized and the RSP performance drops significantly. In these regimes, the OLP can substantially improve the effective dynamics, reaching enhancement factors of up to $P_{R}\approx 2$ for the Santa Fe task and $P_{R}\approx 27$ for the short-term memory task. When the unperturbed dynamics already performs well, the values of $P_{R}$ typically remain close to one, showing that the OLP can match the performance of the unperturbed reservoir. As we discuss in the next analysis based on $P_{R}^{\min}$, identifying operating points where back-action also improves upon the best unperturbed configuration requires a more stringent comparison.
Figure 3: Minimum performance ratio analysis for the forward prediction task (top) and for the short-term memory task (bottom). A $P_{R}^{\min}$ in the range $[0,1)$ shows that the best performing RSP is performing better than the OLP in that point of the ghg-h map. A $P_{R}^{\min}$ around 1 shows that the OLP is performing as good as the best performing RSP, and over 1 indicates an over-performance. The optimal parameters and corresponding $P_{R}^{\min}$ values for each task are: forward prediction — RSP: $h^{\prime*}=0.346$; OLP: $g^{*}=0.355$, $h^{*}=0.273$, $P_{R}^{\min}=1.006$; short-term memory — RSP: $h^{\prime*}=0.084$; OLP $g^{*}=0.256$ , $h^{*}=0.066$ , $P_{R}^{\min}=1.029$.
In Figure 3 we show the behaviour of $P_{R}^{\min}$ across the same $(g,h)$ grid used in the previous colormaps (33 by 40 points, infinite-shot assumption) for the forward prediction task on the Santa Fe dataset (top) and for the short-term memory task (bottom). Since $P_{R}^{\min}$ compares the OLP with the best-performing RSP over all scanned values of $h$, it provides a stricter benchmark.
For both tasks, a region with $P_{R}^{\min}>1$ clearly emerges, showing that back-action plays a role even for those OLP configurations that yield the highest capacities. The effect is smaller than in the $P_{R}$ maps, but remains visible: for the Santa Fe task the advantage concentrates at small $h$, where the unperturbed dynamics is most limited, while for the short-term memory task a narrower enhancement region appears around $h\sim 10^{-1}$, consistent with the location of the optimal unperturbed performance.
Overall, the two sets of colormaps convey a consistent picture. The $P_{R}$ maps show that back-action can substantially enhance the performance of the reservoir, especially in regions where the unperturbed dynamics is most limited. In these regimes the OLP not only compensates for the information extracted during the indirect measurement, but can even surpass the corresponding unperturbed dynamics. The $P_{R}^{\min}$ maps confirm that this positive contribution persists also for the best-performing OLP configurations, although the effect is naturally smaller due to the stricter benchmark. Together, these observations demonstrate that a properly tuned indirect-measurement protocol can retain the full capabilities of the unperturbed dynamics while also providing genuine performance gains in suitable regions of parameter space.
@@
To stress this last point, in Figure 4, we compare the different sum capacities of the QRC running with the RSP and with the OLP in the case in which $g$ is fixed to a non-optimal value and in the case in which it is fixed to the value that maximized $P_{R}^{\min}$ in Figure 3. As a first observation, we can refer to [martinez2021dynamical] to justify the dependence of the performance on the $h$ value. In fact, along the horizontal axis, the quantum system undergoes a transition from a localized regime to an ergodic one and it was showcased that being in the ergodic regime is a preferential condition to perform reservoir computing. Moreover, what emerges clearly is that optimizing the measurement strength, rather than simply minimizing back-action, is crucial for improving the protocol.
@@
Figure 4: Sum capacity comparison of the RSP with respect to the OLP with and without optimizing the strength of the measurement $g$. With $g^{*}$ we denote the $g$ value that maximizes the $P_{R}^{\min}$. The top panel shows the results for the forward prediction task while the bottom panel is for the short-term memory task.
@@
The previous analysis is related to the sum capacity of the QRC across two different tasks. However, for different sub-tasks, defined byăvarying $\eta$, the performance may systematically differ. In this context, Figure 5 compares two points in the ghg-h parameter space explored in Fig. 3, the one that maximized $P_{R}^{\min}$ and a non-optimal point. We then evaluate the performance for each different sub-task relative to the performance of the RSP at optimal $h$.